OpenStudy (anonymous):

simplify √(1-sinθ)(1+sinθ) ±sin θ cos θ ±tan θ √sine θ

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@agent0smith

OpenStudy (mrnood):

first of all simplify (a+b)(a-b) = a^2-b^2 then note that cos^2 - 1-sin^2 then simplify again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I have know idea how to solve this

OpenStudy (mrnood):

first of all simplify (a+b)(a-b) = a^2-b^2 correction from above: then note that cos^2 = 1-sin^2 then simplify again

OpenStudy (anonymous):

How would I simplify

OpenStudy (mrnood):